OpenStudy (anonymous):

Evaluate the given function..............( see the comments for the question)

OpenStudy (anonymous):

here is the question

OpenStudy (anonymous):

Please explain how did you arrive at the answer, I don't know how to use L'Hopital's rule, nor has my professor taught us that.

OpenStudy (anonymous):

so , I found 1/10

OpenStudy (anonymous):

Lemme try.....

OpenStudy (anonymous):

Sorry, its the wrong answer

OpenStudy (anonymous):

ok let me check my work .)

OpenStudy (anonymous):

r u sure it's wrong .s

OpenStudy (anonymous):

yeah, i did check it, the question system says, its wrong.

OpenStudy (anonymous):

Okay, I double checked, your answer is right, sorry...

OpenStudy (anonymous):

How did you get it??

OpenStudy (anonymous):

Change y=1/x and y-->0 sqrt(25/y^2 + 1/y) - 5/y= sqrt[(25 + y)/y^2] - 5/y=[sqrt(25+y) - 5]/y Let's complete the square from 25+y by +/-(y/10)^2 25+y=25 + 2* 5* y/10 + (y/10)^2 - (y/10)^2=(5+y/10)^2-(y/10)^2 Continue sqrt[(5+y/10-5]/y=1/10 lim(x-->infinity)=1/10

OpenStudy (anonymous):

okay why did we do the first step?

OpenStudy (anonymous):

@Emineyy, why did u do the first step??

OpenStudy (anonymous):

Dude, could u tell me how to approach solving this question??

OpenStudy (debbieg):

Wait.... am I missing something? It's: \[\Large \lim_{x \rightarrow \infty}\sqrt{25x^2+x}-5x\] right? The limit of a square root is the square root of a limit. The limit of a polynomial as \({x \rightarrow \infty}\) is determined by the behavior of the leading term (sign of coefficient, and degree). The limit of a difference is the difference of the limits. I think that's all you need to use here??

OpenStudy (anonymous):

yup, you got the question right.

OpenStudy (anonymous):

the answer is correct

OpenStudy (anonymous):

I checked it, initially the website said, its wrong, but when I double checked it came out to be right

OpenStudy (debbieg):

It said 1/10 was the limit?

OpenStudy (anonymous):

yes

OpenStudy (debbieg):

Oh nvr mind... that's what Wolfram says too. OK, I'm bowing out, I'm obviously missing something here, lol.

OpenStudy (anonymous):

:)))

OpenStudy (anonymous):

Hey, you are back, help me out here bud

OpenStudy (anonymous):

So, I don't get why we did the first step and why we did the last step where 5+y/10-5 = 1/10

OpenStudy (anonymous):

@Emineyy

OpenStudy (anonymous):

help me out bud....@Emineyy

OpenStudy (anonymous):

I mean, isn't y = 1/x and if that's so, then shouldn't it be 0??

OpenStudy (debbieg):

Well, I do see why it isn't as simple as I thought initially... duh.... The limit of the square root part is + infinity The limit of the -5x is - infinity So it isn't like they can just be added. <feeling silly> So let me see if I can make sense of her approach... she's making a substitution, re-writing with \(y=1/x\) so then \(x=1/y\) and \(x^2=1/y^2\). \(\sqrt{25/y^2 + 1/y} - 5/y= \sqrt{(25 + y)/y^2} - 5/y=\dfrac{\sqrt{(25+y)} - 5}{y} \) =====I don't understand what the purpose of all this completing the square stuff was about...??===== Let's complete the square from 25+y \(25+y=25 + 2\cdot5\cdot \dfrac{y}{10} + (\dfrac{y}{10})^2 - (\dfrac{y}{10})^2=(5+y/10)^2-(y/10)^2 \) OK, so: \(25+y=(5+y/10)^2-(y/10)^2 \) ===================================== Now back to the limit... with the substitution, we have: \(\lim_{y \rightarrow 0}\dfrac{\sqrt{(25+y)} - 5}{y} \) which is a 0/0 form and L'hopital's rule applies. \(\dfrac{d}{dy}(25+y)^{1/2} - 5=\dfrac{1}{2}(25+y)^{-1/2}=\dfrac{1}{2}\dfrac{1}{(25+y)^{1/2}}\) so: \(\lim_{y \rightarrow 0}\dfrac{1}{2}\dfrac{1}{(25+y)^{1/2}}=\dfrac{1}{2\cdot5}=\dfrac{1}{10}\) (And the derivative of the "y" den'r is just 1).